OpenStudy (jojo12g):

A triangular pyramid has the same base and height as a triangular prism. The volume of the pyramid is 31 cubic millimeters. What is the volume of the triangular prism? mm3

OpenStudy (meowimacat):

Are there answers? Or nah?

OpenStudy (jojo12g):

oh no

OpenStudy (jojo12g):

theres not

OpenStudy (meowimacat):

Whale I dunno, is it multiple choice? Oh.. there aren't? Okay... one sec then :3

OpenStudy (jojo12g):

@Will.H

OpenStudy (meowimacat):

Okay, I am so tired... But I'm sure if they both are the same height and have the same base dimensions, they should be the same volume. To check that, you should use the formula for volume of the pyramid to find the dimensions and apply them for the prism... I was up till 4 am this morning.. I'm no good at explaining right now. I'm sorry..

OpenStudy (jojo12g):

lol take a nap then dude

OpenStudy (welshfella):

volume of pyramis = 1/3 * height * area of base volume of prism = height * area of base

OpenStudy (will.h):

@welshfella is right :)

OpenStudy (welshfella):

Volum pyramid 1/3 * h * A ------------- = ---------- = 1/3 h * A Volume prism

OpenStudy (welshfella):

so if Volume prism = 31 volume of pyramid = ?

OpenStudy (jojo12g):

31?